OpenStudy (anonymous):

use half angle identity to find the exact value of each function: tan(pi/8)

OpenStudy (anonymous):

Use the formula for tan((1/2)x) where x = pi/4 Because 1/2 of pi/4 is pi/8 and we know that tan (pi/4) is a popular angle = 1. Use formula in your textbook. tan (1/2 x) = sqrt ((1-cos x)/(1+cos x))

OpenStudy (anonymous):

Thanks, I should have said this in the beginning, but I can get that far and then I work it and I don't do it right. So I know this is wrong, but I got sqrt (2 + sqrt (2)). Do I multiply the numerator and denominator by 1/2 first or what?

OpenStudy (anonymous):

Isnt it sqrt((1=srt(2)/2 divided by (1 + sqrt (2)/2)? I dont have pencil/paper in front of me to simplify that...but you would want to simplify the numerator and denominator..and the 2's in both top and bottom will cancel in this complex fraction.

OpenStudy (anonymous):

so then it is (1- sqrt2) / (1 + sqrt2)?